Two six sided dice, with sides numbered 1 through 6 are rolled. If the dice are rolled 9000 times, about how many times would you expect the dice to have a sum of 8?
Found 2 solutions by alicealc, amoresroy:
Answer by alicealc(293)

2 dice = 6*6 = 36 possibilities
sum of 8:
(2,6)
(3,5)
(4,4)
(5,3)
(6,2)

probability = 5/36
expectation frequency = 5/36 * 9000 = 45000/36 = 5000/4 = 1250 times

Answer by amoresroy(361)


sum of 8 combinations
2,6
3,5
4,4
5,3
6,2
There 5 combinations out of 36.
Let x = the number of times we exoect the dice to have a sum of 8
x = 9000*5/36
x = 1,250